# Dense 3D Displacement Estimation for Landslide Monitoring via Fusion of TLS Point Clouds and Embedded RGB Images

### v1

@@ 5.2 Accuracy evaluation
Tab. 1 shows the quantitative results of our method for both landslides. We compare the estimated 3D displacements of our method with external observations: TS prisms on the Brienz dataset and GNSS stations on the Mattertal dataset. For clarity of interpretation, the coordinate system for both datasets is defined following the right-hand rule, where the Y-axis approximately corresponds to the horizontal projection of the mean line-of-sight (LOS) direction, the X-axis is perpendicular to the Y-axis horizontally, and the Z-axis points vertically upward. The displacement magnitude deviations (i.e., $|\Delta D_{S}|$) between our estimates and the external observations are 0.15 m and 0.25 m, respectively. They are lower than the corresponding average footprint sizes of 0.23 m and 0.68 m, respectively. Additionally, we manually select corresponding points at the second epoch (i.e., target epoch) to derive displacements that accurately reflect the input data used in both our method and other baseline approaches. The deviations between our method and the manually selected reference displacements are 0.07 m and 0.20 m for the Brienz and Mattertal datasets, respectively. These deviations are smaller than the average scan resolutions of 0.08 m and 0.30 m for these two datasets, respectively. This demonstrates the high accuracy of the displacement estimation by our method. Furthermore, our method has larger deviations with external observations than with manual reference data, which can be attributed to georeferencing errors. To avoid this, we use the manually picked points as the reference for comparing our method with baseline methods.
@@
Ours (m) $|\mathrm{Ours-External}|$ (m) $|\mathrm{Ours-Manual}|$ (m)
@@
Average 4.51 0.68  -  -  - - 0.37 0.05 0.10 0.25 0.28 0.05 0.19 0.20
Figure 8: Quantitative comparison of estimated displacements from different methods w.r.t. the manual reference, focusing on deviations along the X- and Y-axis directions (i.e., $\Delta D_{X}$ and $\Delta D_{Y}$). Dashed circles and grid lines are added to improve readability. Similar deviation patterns are observed along the X- and Z-axis directions as well (see B).
We further compare the estimated displacements of our method with those of baseline methods in Fig. 8, where we use manually picked points as the reference and compare the deviations between the different methods and the reference. Our method shows maximum deviations of 0.17 m along the X-axis for the Brienz dataset and 0.48 m along the Y-axis for the Mattertal dataset. Among all baseline methods, F2S3 achieves the closest performance to ours, with maximum deviations of 0.18 m along the Y-axis for the Brienz dataset and 0.57 m along the Y-axis for the Mattertal dataset. It is worth noting that all baseline methods fail to provide valid estimates corresponding to all external observations (e.g., Piecewise ICP only provides two valid estimates for all four GNSS stations on the Mattertal dataset) due to their reliance on only one type of information (either 3D point cloud or 2D RGB information). In contrast, our method provides estimates around all external observations. Piecewise ICP (Friedli and Wieser, 2016) exhibits the largest deviations on both datasets, primarily due to its approach of uniformly partitioning and subsequently matching the closest patches based on Euclidean distance. The RGB-Guided approach shows large deviations along the Y-axis direction. An investigation in Wang et al. (2023) suggests that, even though the manufacturer provides co-registration between the TLS scanner and its built-in camera(s), misalignment errors of several pixels can still occur between the point clouds and the images. These errors persist across epochs due to variations in the camera poses associated with the captured images, even when the same scanner is used. This is particularly problematic on the Mattertal dataset, where the average GSD is 0.2 m, resulting in potential meter-level displacement errors. In our approach, 2D information mainly contributes to patch matching. Due to the refinement, our method with fused 3D and 2D information still is comparable in accuracy to the 3D-based approach, F2S3.
5.3 Spatial coverage evaluation
We present the 3D displacements estimated by our method in Fig. 9 and Fig. 10 for the Brienz and Mattertal datasets, respectively. These results demonstrate that our method can estimate 3D displacements across most landslide areas, with spatial coverage of 79% and 97% w.r.t. the source point cloud for the Brienz and Mattertal datasets, respectively. The Brienz landslide exhibits significant variation in displacement magnitude across the area, whereas the Mattertal landslide remains relatively stable, with only a few localized areas showing notable movement between the observed epochs. To further analyze the displacement characteristics, we select a ROI from the highly active areas of each landslide, where several external observations are also included in the corresponding ROI. In the Brienz ROI (see Fig. 9), the 3D displacement vectors estimated by our method have a mean magnitude of approximately 2.5 m and mean azimuth and elevation angles of approximately 135° and -45°, respectively. This suggests displacement toward the southeast. In contrast, most of the displacement magnitudes in the Mattertal landslide ROI (see Fig. 10) are close to zero, and the azimuth and elevation angles exhibit large spreads without consistent direction trends. This indicates a lack of coherent surface motion during the observed period.
Figure 9: Estimated 3D displacements on the Brienz landslide. (a): Displacement magnitudes across the entire landslide (ROI highlighted in  red). (b) Azimuth direction of displacements within the ROI. (c): Elevation direction of displacements within the ROI. The source point cloud is shown in gray as the background for better readability, which applies to all other figures in this section.
Figure 10: Estimated 3D displacements on the Mattertal landslide. (a): Displacement magnitudes across the entire landslide (ROI highlighted in  red). (b) Azimuth direction of displacements within the ROI. (c): Elevation direction of displacements within the ROI.
Fig. 11 and Fig. 12 present qualitative comparisons between our method and baseline methods for the ROIs of the Brienz and Mattertal datasets, respectively. While Piecewise ICP achieves high spatial coverage (93% and 95%) on both ROIs, it underestimates the actual displacements on the Brienz dataset and produces outlier estimates on the Mattertal dataset. In contrast, the other three methods produce results that better reflect the actual displacement trends. For both ROIs, F2S3 achieves spatial coverage of 38% and 52%, and the RGB-Guided approach achieves spatial coverage of 45% and 66%. These are both much lower than the coverage of 82% and 90% achieved by our method. This is attributed to the partition-based coarse-to-fine estimation and the integrated use of both 3D and 2D information in our method. Furthermore, the RGB-Guided approach produces many noisy estimates on the Mattertal dataset, primarily due to the aforementioned co-registration misalignment errors (see Sec. 5.2). In our method, noisy correspondences from the 2D source are largely suppressed due to the adaptive balance between 3D and 2D information.
Figure 11: Qualitative comparison of estimated displacement magnitudes using different methods on the ROI of the Brienz landslide.
Figure 12: Qualitative comparison of estimated displacement magnitudes using different methods on the ROI of the Mattertal landslide.
5.4 Effectiveness of the fusion module
We evaluate the effectiveness of the fusion module, with qualitative results shown in Fig. 13. For comparison, we define two ablated variants: Ours-2D, which performs both coarse and fine matching using only RGB image information, and Ours-3D, which relies solely on geometric information from the point cloud. We observe that the RGB-Guided method overestimates the displacement magnitudes, consistent with quantitative results in Sec. 5.2. This may be attributed to the co-registration error, which also decreases the quality of 3D displacements estimated by Ours-2D. Conversely, Ours-3D yields more accurate estimates but fails in regions dominated by debris, where geometric features are less discriminative. By integrating both 2D RGB and 3D geometry through the fusion module, our method improves spatial coverage to 90% compared to 87% of Ours-3D, while mitigating noisy correspondences introduced by 2D RGB matching alone.
Figure 13: Comparison of our method using only 2D information, only 3D information, and both 2D and 3D information on the ROI of the Mattertal landslide.
5.5 Impact of multiple levels of partitions
We analyze the impact of using different partition strategies, with qualitative results shown in Fig. 14. We apply each of our three hierarchical levels (level i, ii, iii) of partitions independently using the same matching pipeline. The supervoxel-based partition adopted in F2S3 serves as a baseline and represents a single-level approach. It produces displacements with 74% spatial coverage, which is higher than our level i (55%) but lower than level ii (79%). While level iii offers higher spatial coverage (90%), it tends to over-smooth displacements by breaking object boundaries. In contrast, our proposed hierarchical method (cf. ours) integrates all three levels incrementally, yielding 90% coverage while better preserving local structural details.
Figure 14: Comparison of our method using different partitioning strategies or levels on the ROI of the Mattertal landslide.
5.6 Effectiveness of the refinement module
We analyze the impact of the refinement module and present the qualitative results in Fig. 15. We compare three cases: (i) The original 3D result obtained directly from the pre-trained deep learning model; (ii) the result obtained using our method without the refinement module; and (iii) the result obtained using our method with the refinement module. The original 3D result has relatively low spatial coverage (54%) and many noisy estimates, which can be attributed to two factors: 1) The 3D matching fails to produce a sufficient number of matches due to the limitations of geometric features in some regions, where RGB features could contribute, and 2) the absence of the refinement module. When we compare results with and without the refinement module, we find that our method without refinement yields an output with higher spatial coverage (92%) but more noisy estimates. In contrast, our method with the refinement module effectively eliminates low-quality patch matches, primarily consisting of noisy correspondences, by applying a strict and practical filtering criterion. This improves overall match reliability while maintaining high spatial coverage (90%) of the output.
Figure 15: Comparison of estimated displacement magnitudes: original 3D output, our method without the refinement module (cf. Ours w/o refine.), and our method with the refinement module (cf. Ours) on the ROI of the Mattertal landslide.
5.7 Limitations and potential improvements
Limitations: Our approach has one distinct limitation: it relies on feature distinctiveness for accurate point-wise or pixel-wise matching. In regions with repetitive geometric structures or weak radiometric texture, e.g., debris-covered surfaces, or in cases of severe occlusion and topographic disruption caused by large toppling or rotational failures, the method may fail to establish accurate correspondences, resulting in missing displacement estimates. Potential improvements: (i) We adopt the 3D patch-based model from (Poiesi and Boscaini, 2021) and the image matching algorithm from (Wang et al., 2024), both of which are commonly used in related applications. While they currently provide strong performance, we believe that future advances in feature learning and image matching may further improve the spatial coverage and accuracy. (ii) The method is applied to TLS datasets with images captured by built-in, calibrated cameras. Since these cameras are closely aligned with the scanner in terms of viewpoint, fusion benefits are limited to the feature level. We anticipate that integrating images from different viewpoints, such as stationary cameras or UAV platforms, could introduce complementary information at the data level. We leave this exploration for future work.
6 Conclusion
This paper presents a hierarchical partition-based approach for estimating dense 3D DVF in TLS-based landslide monitoring. The method optimizes the matching pipeline in a coarse-to-fine manner using both 3D point clouds and 2D RGB images. By integrating multi-level partition information, we effectively capture structural variations and mitigate the occurrence of noisy estimates while maintaining high spatial coverage. Furthermore, rather than requiring annotated landslide-specific training datasets, which are hard to obtain, the method leverages features extracted from models pre-trained on publicly available indoor datasets. Experiments on two real-world landslide datasets with different displacement patterns validate the effectiveness of the proposed method: it produces displacement estimates with higher spatial coverage than the current state-of-the-art method (F2S3) while maintaining comparable accuracy. Although this method has been verified for TLS-based landslide monitoring, it can be extended to other types of point clouds and other monitoring applications. Future research will focus on exploring the integration of TLS point clouds with RGB images captured from external cameras, such as those mounted on UAVs. This could provide multi-perspective data, enabling landslide monitoring with an enhanced spatial level and coverage.
CRediT authorship contribution statement
Zhaoyi Wang: Conceptualization, Methodology, Data curation, Investigation, Formal analysis, Validation, Visualization, Writing – original draft, Writing – review & editing, Software. Jemil Avers Butt: Conceptualization, Methodology, Formal analysis, Writing – review & editing, Supervision. Shengyu Huang: Conceptualization, Methodology. Tomislav Medić: Methodology, Writing - Review & Editing. Andreas Wieser: Conceptualization, Writing – review & editing, Resources, Supervision, Project administration, Funding acquisition.
During the preparation of this work the authors used ChatGPT in order to assist in refining the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Declaration of competing interest
Acknowledgments
The Brienz dataset was provided by Robert Kenner (Institute for Snow and Avalanche Research), and total station observations were provided by Stefan Schneider (CSD INGENIEURE AG). The data collection of the Mattertal dataset was financially supported by the Federal Office for the Environment (FOEN) and by the Swiss Geodetic Commission (SGC). Helena Laasch has shared the code for Piecewise ICP. The first author was supported by the China Scholarship Council (CSC).
Appendix A Comparison with averaged displacements
@@
Following (Gojcic et al., 2021; Shi et al., 2025; Wang et al., 2025a), for each external observation or manually pickled point, we select a set of estimated displacement vectors within a certain radius. We use a 15 m radius for both datasets in default. However, for point T5, we observe a significant deviation in the mean absolute deviation (MAD) due to local non-rigid motion. Therefore, we reduce the radius to 2.5 m for point T5. The average deviations between the estimated displacement magnitudes of our method and the external observations are 0.12 m and 0.30 m for the two datasets, respectively, both smaller than the corresponding footprint sizes of 0.23 m and 0.68 m. The average deviations between estimated displacement magnitudes of our method and external observations are 0.10 m and 0.12 m, both smaller than the average scan resolutions of 0.08 m and 0.30 m. Theses findings are consistent with our analysis based NN search in Section 5.2. The MAD of the estimated displacement magnitudes are 0.15 m and 0.24 m, respectively, suggesting the high precision of the displacement magnitudes estimated by our method.
@@
Ours (m) $|\mathrm{Ours-External}|$ (m) $|\mathrm{Ours-Manual}|$ (m) MAD (m)
@@
Average 4.51 0.68  -  -  - - 0.32 0.08 0.11 0.30 0.15 0.07 0.20 0.12 0.28 0.16 0.14 0.24
Appendix B Deviation patterns along the X- and Z-axis directions

### v2

@@ 5.2 Accuracy evaluation
We compare the estimated 3D displacements of our method with external observations: TS prisms on the Brienz dataset and GNSS stations on the Mattertal dataset. We first present the quantitative results of our method for both landslides in Table 1, where we use NN search to find the estimates of corresponding external observations. Although the estimated displacement magnitudes may appear large, they fall within the range documented for the Swiss Alps in previous studies (see D). For clarity of interpretation, the coordinate system for both datasets is defined following the right-hand rule, where the Y-axis approximately corresponds to the horizontal projection of the mean line-of-sight (LoS) direction, the X-axis is perpendicular to the Y-axis horizontally, and the Z-axis points vertically upward. The displacement magnitude deviations (i.e., $|\Delta D_{S}|$) between our estimates and the external observations are 0.15 m and 0.25 m, respectively. The larger deviation observed on the Mattertal dataset is mainly due to the fact that the mean footprint size is 0.68 m, which is larger than the 0.23 m size on the Brienz dataset. Hence, the 0.25 m deviation falls within the expected accuracy range of the Mattertal dataset. Additionally, we manually select corresponding points at the second epoch (i.e., target epoch) to derive displacements that accurately reflect the input data used in both our method and other baseline approaches. The deviations between our method and the manually selected reference displacements are 0.07 m and 0.20 m for the Brienz and Mattertal datasets, respectively. These deviations are smaller than the mean scan resolutions of 0.08 m and 0.30 m for these two datasets, respectively. This demonstrates the high accuracy of the displacement estimation by our method. Furthermore, our method has slightly larger deviations with external observations than with manual reference data, which can be attributed to georeferencing errors. To avoid this, we use the manually picked points as the reference for comparing our method with baseline methods.
@@
Ours (m) $|\mathrm{Ours-External}|$ (m) $\downarrow$ $|\mathrm{Ours-Manual}|$ (m) $\downarrow$
@@
Mean 4.51 0.68 0.37 0.05 0.10 0.25 0.28 0.05 0.19 0.20
@@
We further present the quantitative results of our method for both landslides in Table 2, where we select a set of estimated displacement vectors within a certain radius and compare their mean with corresponding external observations or manually picked points. This comparison aligns with the evaluation used in previous studies (Gojcic et al., 2021; Shi et al., 2025; Wang et al., 2025a). We use a 15 m radius for both datasets by default. However, for point T5, we observe a significant deviation in the mean absolute deviation (MAD) due to local non-rigid motion. Therefore, we reduce the radius to 2.5 m for point T5. The mean deviations between the estimated displacement magnitudes from our method and the external observations are 0.12 m and 0.30 m for the two datasets, respectively, both smaller than the corresponding footprint sizes of 0.23 m and 0.68 m. The mean deviations between the estimated displacement magnitudes from our method and the manually selected reference displacements are 0.10 m and 0.12 m, both smaller than the mean scan resolutions of 0.08 m and 0.30 m. These findings are consistent with our analysis based on NN search in Section 5.2. The MAD of the estimated displacement magnitudes are 0.15 m and 0.24 m, respectively, suggesting the high precision of the displacement magnitudes estimated by our method.
@@
Ours (m) $|\mathrm{Ours-External}|$ (m) $\downarrow$ $|\mathrm{Ours-Manual}|$ (m) $\downarrow$ MAD (m) $\downarrow$
@@
Mean 4.51 0.68 0.32 0.08 0.11 0.30 0.15 0.07 0.20 0.12 0.28 0.16 0.14 0.24
Figure 8: Quantitative comparison of estimated displacements from different methods w.r.t. the manual reference, focusing on deviations along the X- and Y-axis directions (i.e., $\Delta D_{X}$ and $\Delta D_{Y}$). Dashed circles and grid lines are added to improve readability. Similar deviation patterns are observed along the X- and Z-axis directions as well (see A).
Additionally, we compare the estimated displacements of our method with those of baseline methods in Figure 8, where we use manually picked points as the reference and compare the deviations between the different methods and the reference. Our method shows maximum deviations of 0.17 m along the X-axis for the Brienz dataset and 0.48 m along the X-axis for the Mattertal dataset. Among all baseline methods, F2S3 achieves the closest performance to ours, with maximum deviations of 0.18 m along the Y-axis for the Brienz dataset and 0.57 m along the Y-axis for the Mattertal dataset. All baseline methods fail to provide valid estimates corresponding to all external observations (e.g., Piecewise ICP only provides two valid estimates for all four GNSS stations on the Mattertal dataset) due to their reliance on only one type of information (either 3D point cloud or 2D RGB information). In contrast, our method provides estimates around all external observations. Piecewise ICP (Friedli and Wieser, 2016) exhibits the largest deviations on both datasets, primarily due to its approach of uniformly partitioning and subsequently matching the closest patches based on Euclidean distance. The RGB-guided approach shows large deviations along the Y-axis direction. An investigation in Wang et al. (2023) suggests that, even though the manufacturer provides co-registration between the TLS scanner and its built-in camera(s), misalignment errors of several pixels can still occur between the point clouds and the images. These errors persist across epochs due to variations in the camera poses associated with the captured images, even when the same scanner is used. This is particularly problematic on the Mattertal dataset, where the mean GSD is 0.2 m, resulting in potential meter-level displacement errors. In our approach, 2D information mainly contributes to patch matching. Due to the refinement, our method with fused 3D and 2D information still is comparable in accuracy to the 3D-based approach, F2S3.
5.3 Spatial coverage evaluation
We present the 3D displacements estimated by our method in Figure 9 and Figure 10 for the Brienz and Mattertal datasets, respectively. The results demonstrate that our method can estimate 3D displacements across most landslide areas, with spatial coverage of 79% and 97% w.r.t. the source point cloud for the Brienz and Mattertal datasets, respectively. The Brienz landslide exhibits significant variation in displacement magnitude across the area, whereas the Mattertal landslide remains relatively stable, with only a few localized areas showing notable movement between the observed epochs. To further analyze the displacement characteristics, we select a ROI from the highly active areas of each landslide, where several external observations are also included in the corresponding ROI. In the Brienz ROI (see Figure 9), the 3D displacement vectors estimated by our method have a mean magnitude of approximately 2.5 m and mean azimuth and elevation angles of approximately 135 and -45, respectively. This suggests displacement toward the southeast. In contrast, most of the displacement magnitudes in the Mattertal landslide ROI (see Figure 10) are close to zero, and the azimuth and elevation angles exhibit large spreads without consistent direction trends. This indicates a lack of coherent surface motion during the observed period.
Figure 9: Estimated 3D displacements on the Brienz landslide. (a): Displacement magnitudes across the entire landslide (ROI highlighted in  red). (b): The azimuth direction of displacements within the ROI. (c): The elevation direction of displacements within the ROI.
Figure 10: Estimated 3D displacements on the Mattertal landslide. (a): Displacement magnitudes across the entire landslide (ROI highlighted in red). (b) The azimuth direction of displacements within the ROI. (c): The elevation direction of displacements within the ROI.
Figure 11 and Figure 12 present qualitative comparisons between our method and baseline methods for the ROIs of the Brienz and Mattertal datasets, respectively. While Piecewise ICP achieves high spatial coverage (93% and 95%) on both ROIs, it underestimates the actual displacements on the Brienz dataset and produces outlier estimates on the Mattertal dataset. In contrast, the other three methods produce results that better reflect the actual displacement trends. For both ROIs, F2S3 achieves spatial coverage of 38% and 52%, and the RGB-guided approach achieves spatial coverage of 45% and 66%. These are both much lower than the coverage of 82% and 90% achieved by our method. This is attributed to the hierarchical partitioning-based coarse-to-fine estimation and the integrated use of both 3D and 2D information in our method. Furthermore, the RGB-guided approach produces many noisy estimates on the Mattertal dataset, primarily due to the aforementioned co-registration misalignment errors (see Section 5.2). In our method, noisy correspondences from the 2D source are largely suppressed due to the adaptive balance between 3D and 2D information.
Figure 11: Qualitative comparison of displacement magnitudes estimated by different methods on the ROI of the Brienz landslide. Cov. denotes spatial coverage. The results highlight variations among the methods in both spatial coverage and their ability to capture large displacements.
Figure 12: Qualitative comparison of displacement magnitudes estimated by different methods on the ROI of the Mattertal landslide, where the results produced by different methods differ in both spatial coverage and local displacement patterns.
5.4 Effectiveness of the fusion module
We evaluate the effectiveness of the fusion module, with qualitative results shown in Figure 13. For comparison, we define two ablated variants: Ours-2D, which performs both coarse and fine matching using only RGB image information, and Ours-3D, which relies solely on geometric information from the point cloud. We observe that the RGB-guided method overestimates the displacement magnitudes, consistent with quantitative results in Section 5.2. This can be attributed to the co-registration error, which also decreases the quality of 3D displacements estimated by Ours-2D. Conversely, Ours-3D yields more accurate estimates but fails in regions dominated by debris, where geometric features are less discriminative. Specifically, at station G1, both Ours (fusion) and Ours-3D have 0.07 m deviations of the estimated displacement magnitudes, while Ours-2D fails to provide a valid estimate. At station G2, Ours-2D exhibits a large deviation of 1.01 m, while Ours-3D fails to provide a valid estimate. Ours has only a 0.28 m deviation. These results highlight the ability of our method to mitigate errors from 2D-only matching while retaining the accuracy of 3D information. Furthermore, by integrating both 2D RGB and 3D geometry through the fusion module, our method improves spatial coverage to 90% compared to 87% of Ours-3D.
Figure 13: Comparison of our method using only 2D information, only 3D information, and both 2D and 3D information on the ROI of the Mattertal landslide, the results of ours achieves the highest spatial coverage while maintaining the high accuracy.
5.5 Impact of multiple levels of partitions
We analyze the impact of using different partitioning strategies, with qualitative results shown in Figure 14. We apply each of our three hierarchical levels (level i, ii, iii) of partitions independently using the same matching pipeline. The supervoxel-based partitioning adopted in F2S3 serves as a baseline and represents a single-level approach. It produces displacements with 74% spatial coverage, which is higher than our level i (55%) but lower than level ii (79%). While level iii offers higher spatial coverage (90%), it tends to over-smooth displacements by breaking object boundaries. In contrast, our proposed hierarchical method (cf. ours) integrates all three levels incrementally, yielding 90% spatial coverage while better preserving local structural details.
Figure 14: Comparison of our method using different partitioning strategies and levels on the ROI of the Mattertal landslide. The optimal configuration (ours) is adopted, as it produces the highest spatial coverage in both quantitative and qualitative evaluations.
5.6 Effectiveness of the refinement module
We analyze the impact of the refinement module and present the qualitative results in Figure 15. We compare three cases: (i) The original 3D result obtained directly from the pretrained deep learning model; (ii) the result obtained using our method without the refinement module; and (iii) the result obtained using our method with the refinement module. The original 3D result has relatively low spatial coverage (54%) and many noisy estimates, which can be attributed to two factors: 1) The 3D matching fails to produce a sufficient number of matches due to the limitations of geometric features in some regions, where RGB features could contribute, and 2) the absence of the refinement module. When we compare results with and without the refinement module, we find that our method without refinement yields an output with higher spatial coverage (92%) but more noisy estimates. In contrast, our method with the refinement module effectively eliminates low-quality patch matches, primarily consisting of noisy correspondences, by applying a strict and practical filtering criterion. This improves overall match reliability while maintaining high spatial coverage (90%) of the output.
Figure 15: Comparison of estimated displacement magnitudes: original 3D output, our method without the refinement module (cf. Ours w/o refine.), and our method with the refinement module (cf. Ours), on the ROI of the Mattertal landslide.
5.7 Runtime analysis
The proposed approach has an overall time complexity of $O(N\log N)$ due to the NN search between 3D points in feature space. We use the GPU implementation of Faiss (Johnson et al., 2019) for NN search. We further compare the runtime between different approaches, with one image used for both the RGB-guided baseline and our approach. A pair of tiles from the Brienz ROI region (512k source and 571k target points) is used for this analysis. To make a fair comparison, we evaluate all methods on a single GeoForce RTX 3090 Ti with an AMD Ryzen 7 5800x 8-Core Processor. The runtimes for the RGB-guided approach (2D RGB approach), and F2S3 (3D geometry approach) are 81 s and 133 s, respectively. Our fusion approach requires 221 s when only partition level ii is used, which is comparable to the sum of the RGB-guided and F2S3 runtimes. Using all three partition levels increases the runtime to 259 s, indicating that higher spatial coverage can be achieved with only a moderate runtime increase. The runtime of the Piecewise ICP is only 11 s, which demonstrates its high processing efficiency.
5.8 Limitations and potential improvements
Limitations: Our approach has several limitations. First, it relies on feature distinctiveness for accurate point-wise or pixel-wise matching. In regions with repetitive geometric structures and weak radiometric textures, e.g., debris-covered surfaces and vegetation regions, or in cases of severe occlusion and topographic disruption caused by large toppling or rotational failures, the method may fail to establish accurate correspondences, resulting in missing displacement estimates. Second, the current fusion strategy performs coarse matching using 3D point features and 2D pixel features independently, without exploiting deep feature-level fusion. This limitation is partly attributed to the scarcity of annotated, landslide-specific training data. Potential improvements: (i) We adopt the 3D patch-based model from (Poiesi and Boscaini, 2021) and the image matching algorithm from (Wang et al., 2024), both of which are commonly used in related applications. While they currently provide strong performance, we believe that future advances in feature learning and image matching may further improve the spatial coverage and accuracy. Fine-tuning the pretrained model with simulated data may be another option. (ii) Due to the lack of dense ground-truth data for evaluation, one could investigate and evaluate the approach using simulation-based datasets. (iii) The method is applied to TLS datasets with images captured by built-in, calibrated cameras. Since these cameras are closely aligned with the scanner in terms of viewpoint, fusion benefits are limited to the feature level. We anticipate that integrating images from different viewpoints, such as stationary cameras (Neyer, 2016) or UAV platforms (Sestras et al., 2025), could introduce complementary information at the data level. We leave this exploration for future work.
6 Conclusion
This paper presents a hierarchical partitioning-based approach for estimating dense 3D DVF in TLS-based landslide monitoring. The method optimizes the matching pipeline in a coarse-to-fine manner using both 3D point clouds and 2D RGB images. By integrating multi-level partition information, we effectively capture structural variations and mitigate the occurrence of noisy estimates while maintaining high spatial coverage. Furthermore, rather than requiring annotated landslide-specific training datasets, which are difficult to obtain, the method leverages features extracted from models pretrained on publicly available indoor datasets. Experiments on two real-world landslide datasets with different displacement patterns validate the effectiveness of the proposed method: it produces displacement estimates with higher spatial coverage than the current state-of-the-art method (F2S3) while maintaining comparable accuracy. Although this method has been verified for TLS-based landslide monitoring, it can be extended to other types of point clouds and monitoring applications. Future research will focus on exploring the integration of TLS point clouds with RGB images captured from external cameras, such as those mounted on UAVs. This could provide multi-perspective data, enabling landslide monitoring with an enhanced spatial resolution and coverage.
CRediT authorship contribution statement
Zhaoyi Wang: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Jemil Avers Butt: Writing – review & editing, Methodology, Formal analysis, Conceptualization. Shengyu Huang: Methodology, Conceptualization. Tomislav Medić: Writing – review & editing, Data curation, Methodology. Andreas Wieser: Writing – review & editing, Supervision, Resources, Project administration, Funding acquisition, Conceptualization.
During the preparation of this work, the authors used ChatGPT in order to assist in refining the readability and language of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Declaration of competing interest
Acknowledgments
We thank the anonymous reviewers for their constructive comments and suggestions, which helped improve the clarity and quality of this manuscript. The Brienz dataset was provided by Robert Kenner (Institute for Snow and Avalanche Research), and total station observations were provided by Stefan Schneider (CSD INGENIEURE AG). The data collection of the Mattertal dataset was financially supported by the Federal Office for the Environment (FOEN, grant 20.0011.PJ /6DDBCAD6D) and by the Swiss Geodetic Commission (SGC). Helena Laasch has shared the code for the Piecewise ICP. The first author was supported by the China Scholarship Council (CSC).
Data and code availability
Our example data and source code are publicly available at https://github.com/gseg-ethz/fusion4landslide.
Appendix A Deviation patterns along the X- and Z-axis directions
See Figure 16.